what is the easiest way to subtract 27 1/2 liters from 18 5/8?
Answer by DWL(56)

The problem here is to subtract 27 1/2 and 18 5/8
These two fractions do not have the same denominators (lower numbers), so we must first find a common denominator of the two fractions, before subtracting them.

For the denominators here, the 2 and 8, a common denominator for both is 8.
With the common denominator
So now our subtraction problem becomes this... 27 4/8 - 18 5/8
These two fractions have the same denominators, but numerator of the first fraction, the 4, is less than the numerator of the second fraction, the 5. This means, before subtracting the fractions, we have to "borrow" a 1 from the whole number 27, making it a 26. We can add this borrowed 1 to the fraction 4/8, like this:
4/8 + 1 is 4/8 + 8/8 = 12/8
After doing all of this, the becomes a , and we can now perform the subtraction...
The problem here is to subtract and 26 12/8 - 18 5/8
Since these two fractions have the same denominators (the numbers under the fraction bar), we can subtract them by simply subtracting the numerators (the 12 - 5 = 7), while keeping the same denominator (the 8).

And, we just subtract the whole number parts, or 26 - 18 = 8.

The final answer is: 8 7/8